In a survey, 25 people were asked how much they spent on their child's last birthday gift. The results were roughly bell-shaped with a mean of $36 and standard deviation of $16. Estimate how much a typical parent would spend on their child's birthday gift (use a 95% confidence level).

Give your answers to one decimal place. Provide the point estimate and margin or error.
Answer #1

Ans:

n=25

df=25-1=24

critical t value=tinv(0.05,24)=2.064

Margin of error=2.064*16/sqrt(25)=6.604

Confidence interval for mean

=36+/-6.604

=(29.396, 42.604)